What is the probability of NOT getting 1 or 2 on a 6 sided dice when rolled 15 times?
P(1or2)=2/6=1/3
P(4<=x<=6)=2/3
when rolled 15 times
(2/3)^15=.0023
How can probability be used to make predictions?
Well... with what I learned from Mrs. Franks, mt math teacher, she said for weather. For example there with be a probability of 75 degrees today.
If the probability of an event is 0.02, then the probability of two such events occurring is 0.022 or 0.0004.
There are 13 hearts in a 52 card deck so the probability of drawing a heart is 13/52 or 1/4.
What is the value of 4800 if the probability is 20?
The value of 4800 is 4800 and probability has nothing to do with it.
What is the probability of spining a 147 or 10?
The answer depends on what it is that you are spinning - or spining (whatever that might be!).
The answer depends on what it is that you are spinning - or spining (whatever that might be!).
The answer depends on what it is that you are spinning - or spining (whatever that might be!).
The answer depends on what it is that you are spinning - or spining (whatever that might be!).
What is the probability that if a month is selected from a calendar it has more than 28 days?
The probability is approx 15/16.
Is a Person who has been sick and in the hospital after illness is probably feeble?
The answer is yes.
Does the male or female pass colorblindess to the offspring?
It is on the X Chromosome (male have XY chromosome Females have XX Chromosome)
What function does the submucosa have?
The submucosa supports the mucosa and joins the mucosa to the bulk of overlying smooth muscle. Blood vessels, lymphatic vessels, and nerves all run through here.
1/1200 =0.0008 < 1%
Why normal distribution is better than other distributions?
It may not be better, but there is a lot of information on the normal distribution. It is one of the most widely used in statistics.
What is the mode in the data sequance 30 10 40 20?
The mode in a set of data is the number which appears most frequently. For instance in the data set 2,2,3,5,6,7,7,7,7,8,9,9, the mode is clearly 7 as it appears most frequently. If there are two numbers that appear equally frequently, such as in the set 1,2,2,3,4,4,4,4,5,6,6,7,8,9,9,9,9, we would say the data is bimodal, and has modes of 4 and 9. If there are 3 equally frequent values, then the data is trimodal and so on. In your example all numbers are equally frequent (each occurring once). In this case we would say there is no mode.
What are the disadvantages of probability sampling?
The three main disadvantages are non-response, representation and expense.
Under probability sampling once the sample is selected the researcher may not interfere with it sample. Specifically, if a large proportion of the sample do not respond, the study will be seriously flawed. With quota sampling, for example, the sampling will continue until the required number of responses are obtained.
The representation issue arises if a small sample is taken from a large population which is made of of sub-populations. Elements within each sub-population are similar to one another but there are important differences between members of different sub-populations. Some sub-populations may end up not being represented in the sample. A typical school-level example is sampling all employees of a company: 2 managers, 90 workers, 8 office staff. If using 10% samples, most (96% of them) will exclude any managers. Stratified sampling will get around that.
The expenses issue can be best illustrated by an example. Suppose you want to sample 1% of all schools in your country. You get a list of all schools in the country and select a 1% random sample. These will be scattered all over the country and you (or your team) will spend a huge amount of time and money travelling from one location to another. A better alternative is to divide up the country into a number of regions, each with [approximately] the same number of schools. Select a 1% sample of the regions [not schools] and then conduct a census of all schools within the selected regions. A variant might be to sample 2% of the regions and, within each region sample 50% of the schools so as to give a 2%*50% = 1% sample overall.
What is the event space for rolling the sum of 8 with two dice?
There are 36 permutations of rolling two dice. Of these, there are five that add up to a sum of 8...
6+2
5+3
4+4
3+5
2+6
This translates to an event (or sample) space of {62, 53, 44, 35, 26}.
The probability, then, of rolling a sum of 8 on two dice is 5 in 36, or about 0.1389.
What is the chance of drawing a 6 form a deck of cards?
There are four 6s in a standard deck of 52 cards, so the chance of drawing a 6 is:
4/52 = 1/13 = 0.0769
What is the mean absolute deviation of this set 3 9 27 81?
Each number is 3 times greater than the previous number
it would be red because the probability is 5/9
If a code contains 4 letters and 2 numbers how many possible outcomes are there?
If the numbers and letters can be repeated then there are 45,697,600 possible outcomes. If the letters and numbers can not be repeated there are 32,292,000 possible outcomes.
Two dice are rolled Find the odds that the score on the dice is either 6 or 12?
There are 36 different number combinationsfor two dice (A & B).
A 12 can only be scored with a 6 on A and a 6 on B
A 6 can be scored with the following combinations:
A = 1, B = 5
A = 2, B = 4
A = 3, B = 3
A = 4, B = 2
A = 5, B = 1
Thus there are 5 ways to score 6 with two dice
Thus there are a total of 6 ways to score 6 or 12 with two dice out of 36 combinations, thus the odds are 6/36 or 1/6